IPP Progress Report for the week 2012-06-25 - 2012-07-01

(Up to IPP Progress Reports)

Eugene Magnier

  • generated a new tag with psphot set to use CONSTANT weight by default; needed to address build issues stemming from the upgrades to support Ubuntu 12.04 / gcc 4.6.3 / Perl 5.14.
  • generating some statistics on the 3pi coverage based on the big 3pi dvo database - I need some figures for the photometry reference ladder paper.
  • ran some tests with PSPS to compare with the DVO version of the SAS database. uncovered a bug in the ipptopsps translation related to the cps sequence of mean magnitude filters (fixed)
  • re-extracted the high-quality sample for the ladder paper based on comments from Eddie. worked on paper updates.

Serge Chastel

  • Sick on Monday
  • Helped Gavin fixing the sticky bit issue for /usr/bin/crontab
  • ippMonitor now displays the stsci storage nodes. However, it seems that there is a bug on the gnuplot version on the cluster (4.2p5). Weird patterns are shown in histogram. I tried with another version 4.2p6 and it was fine.
  • libmysqlclient installation issues (see wiki page)
  • Installed gnuplot 4.6p0 to fix the rendering problem with 4.2p5. Modified czartool plotter accordingly.

Heather Flewelling

  • jtmd sciencey stuff for 1.5 days
  • liason stuff (for WB and LS)
  • czar tuesday/wed
  • attended 8 (!) meetings...
  • investigations of ipptopsps filter problem of objects- testing of new code (in the process learning some python/stilts)
    • found bug in stilts relating to fits column with type = character. You can read them in, you can write to stdout. you can't write to sql, html, latex etc (character gets translated to .)
    • tested gene's filter fix in ipptopsps: verified that the mean mags are consistent and not insane looking (I took ippObjID= 18865000001999 and checked that it looks bad in PSPS, looks good in the new batch, and has the same values in the DVO and the batch)
  • what dvodbs need to be re-relphoted (stacks)
    • MD (current one)
    • LAP.ThreePi?.20110809.V2
    • SAS.20120510
  • transfer of ssp files to ipp004 restarted (current through end of 20120625)

Mark Huber

Bill Sweeney

  • vacation

Chris Waters

  • Confirm that the new mask code is in the trunk, and that the new masks are ready for use.
  • Sent Eddie Schlafly flat field information so he can tie ubercal to JTcal.
  • Set up stsci machines in nebulous, and make the changes to the mhpcc.hosts file to store all warps there. Disabled warp cleanup in LAP and nightly science to allow this archive to fill.
  • Fought noisemaps and SVN-tags, attempting to reproduce the false positive rate we saw before. This is still ongoing.